OpenStudy (anonymous):

Sin theta= -2/4 Find all six trigonometric figures.

OpenStudy (agent0smith):

What does it mean by trigonometric figures? ratios?

OpenStudy (agent0smith):

\[\Large \sin \theta = -\frac{ 1 }{ 2 }\] which we can use to find theta. Sine is negative in the third and fourth quadrants so the angle lies there...

OpenStudy (anonymous):

Or do you mean figures as in like you want to find cos, tan, and then inverse of all of those?

OpenStudy (anonymous):

Yes find tan, cos, csc,cot and sec

OpenStudy (anonymous):

Okay well the easiest way is to just draw the triangle, label your sides (be sure to use a^2+b^2=c^2 to solve for the third side), and then using, SOHCAHTOA, find the trig figures! :)

OpenStudy (agent0smith):

@kaylie94 exactly. But we don't know WHICH quadrant the angle lies in, so we can't know for sure if the other functions are positive or negative :/ was there no other info?

OpenStudy (anonymous):

Hmm this is true. The hypotenuse should be the biggest side (therefore the biggest number), no?

OpenStudy (anonymous):